The raw data has been organised in the frequency table below.





games frequency





6 2





7 5





8 n





9 4





10 4





11 2





12 2





13 2





(a) Write down the value of n.





(b) Calculate the mean number of games played per set.





(c) What percentage of the sets had more than 10 games?





(d) What is the modal number of games?

1
Expert's answer
2022-01-24T16:48:26-0500

a) Taking the value of n=3


games(x) frequency(f) f(x)


6 2 12


7 5 35


8 n=3 24


9 4 36


10 4 40


11 2 22


12 2 24


13 2 26


"b)Mean =\frac{ \sum f(x)}{\sum f}\\
=\frac{219}{24}=9.125\\
c)\frac{6}{24}\times 100
=25\% \\
\\
d)\text{modal number is the number with the highest frequency, so the modal number of the game is 7}"